Find the center of mass of the solid generated by the area bounded by x = 1, x = 3, y = 0

and y = x^2 by revolving about the x-axis.


1
Expert's answer
2022-05-30T23:24:01-0400

When a curve "y=f(x)" is rotated about the "x" - axis, the centre of mass of the solid generated will lie on the "x" – axis because of the symmetry of the curve.


"\bar{x}=\dfrac{\displaystyle\int_{1}^{3}xy^2dx}{\displaystyle\int_{1}^{3}y^2dx}=\dfrac{\displaystyle\int_{1}^{3}x^5dx}{\displaystyle\int_{1}^{3}x^4dx}"

"=\dfrac{[\dfrac{x^6}{6}]\begin{matrix}
 3 \\
 1
\end{matrix}}{[\dfrac{x^5}{5}]\begin{matrix}
 3 \\
 1
\end{matrix}}=\dfrac{910}{363}"

"(\dfrac{910}{363}, 0)"